# On the Fourier transform of the hyperbola and its role in hyperbolic photonics

### v1

@@ I Introduction
In modern material science and photonics, hyperbolas have been playing an important role. When materials (or metamaterials) display extremely anisotropic responses, with oppositely-signed real parts of permittivity components for two orthogonal directions, they lead to a hyperbolic dispersion of the iso-frequency contours of their eigenmodes [9, 10, 11, 12, 13, 14, 15], either in their bulk or at their interface with air. This leads to exciting opportunities for nanoscience, with directional and highly confined optical beams supported by such hyperbolic materials. Their features can be accessed by exciting hyperbolic media with localized sources. Generally, localized excitations of a medium access all available, momenta, leading to an emission pattern that is the spatial Fourier transform of the iso-frequency dispersion contour. For isotropic media, a localized emitter generates a field profile that follows a bullseye pattern, which is indeed the Fourier transform of a circular iso-frequency dispersion curve [16]. Similarly, if a material supports a hyperbolic iso-frequency contour at a given frequency, the emission from a localized emitter will feature a field distribution associated with the Fourier transform of a hyperbola. In order to image its pattern, we can rely on near-field scanning optical experiments, as done in many recent papers [17, 18, 19, 20], or rely on another platform to realize optically the Fourier transform of a spatial pattern, i.e., using an optical lens [21]. Figure 1a shows the experimental image plane distribution created by a lens, while imaging two slits with hyperbolic profile in the object plane. In the following, we analytically evaluate this field pattern, and apply this result to provide physical insights into wave propagation in hyperbolic media and other phenomena featuring the Fourier transform of a hyperbolic curve.
@@ III Graphical representation
We start with this simplest region, which is characterized by a pair of straight lines given by $y=\pm\sqrt{-\nicefrac{{\epsilon_{y}}}{{\epsilon_{x}}}}\,x$. The functional value of the transform for all points in this region is unbounded. This “blow-up” is acceptable since our hyperbola has an infinite support. Note that, as the two asymptotes of the given hyperbola in reciprocal space are described by $k_{y}=\pm\sqrt{-\nicefrac{{\epsilon_{x}}}{{\epsilon_{y}}}}\,k_{x}$ (see Fig. 2a which depicts the case of $\epsilon_{x}<0$ and $\epsilon_{y}>0$), each of the asymptotes is normal to one of the separatrix lines, as shown in Fig. 2b. This orthogonality relation can be understood in the limit of $\kappa\rightarrow 0$ where our hyperbola evolves into the degenerate case of a pair of straight lines. In this limiting case, the corresponding transform (see Eq. (2)) becomes zero in the major and minor regions, and blows up in the separatrix, giving us a pair of straight lines orthogonal to the original degenerate pair, which is expected based on our discussion at the end of Section II. Moreover, this $\kappa\rightarrow 0$ picture also explains why the description of the separatrix is unaltered if the signs of $\epsilon_{x}$ and $\epsilon_{y}$ are reversed. In fact, as the description does not depend on the value of $\kappa$ at all, and hence where the two foci of the hyperbola are located, one can conclude that the separatrix captures the Fourier transform information of the asymptotic behavior of a hyperbola.
@@
Combining the individual picture-pieces of all three regions gives us the complete graphical representation of the Fourier transform of a hyperbola, akin to Fig. 1b, corresponding to the fields radiated by a point source in a hyperbolic medium. Note the coherence among the three regions with a blow-up in the separatrix and decay on either side. Interestingly, the decay in the major region is much slower than that in the minor region. This is due to the fact that, away from the origin, the transform value in the major region scales as $1/{\sqrt{\kappa\Lambda}}$, whereas in the minor region it scales as ${e^{-\kappa\Lambda}}/{\sqrt{\kappa\Lambda}}$ [23]. Now, why is the Fourier transform of a hyperbola the way it is, why does it have fringes on one region while in the other region it is structureless, how do we interpret it physically – we answer these questions in the next section.
@@ IV Physical Interpretation
This additional correspondence gives us the direction of group velocity of the wave for each sector of the field profile (see Figs. 1a and 1b), and affirms that the energy indeed propagates away from the source in a radially outward manner – a fact not obvious for a field distribution that is not circularly symmetric as a bullseye pattern.
@@ V Huygens’ principle in hyperbolic media
Figure 10a also illustrates how Huygens’ principle applied to hyperbolic wavelets naturally leads to the concept of negative refraction [38, 39, 40]. For its key importance in modern optics, we highlight this effect separately in Fig. 11. When light refracts into a hyperbolic medium from an isotropic dielectric through an interface along which the permittivity component of the hyperbolic material is positive, Huygens’ construction, as shown in Fig. 11a, allows light to refract with a negative angle, and hence, energy flows the “wrong” way along the tangential direction. This is to be contrasted with the conventional (positive) refraction picture in Fig. 11b involving two dielectrics with different refractive indices which was originally studied by Huygens [28] through his circular wavelets.
@@ VII Conclusion
The authors thank Viktoriia Rutckaia and J. Ryan Nolen for providing the images used in Figs. 1a and 13a, respectively, and acknowledge support for this work from the Simons Foundation.

### v2

# arXiv:2512.02233v2[physics.optics] 26 Feb 2026

@@ I Introduction
In modern material science and photonics, hyperbolas have been playing an important role. When electromagnetic materials (or metamaterials) display extremely anisotropic responses, with oppositely-signed real parts of permittivity components for two orthogonal directions, they lead to a hyperbolic dispersion of the iso-frequency contours of their eigenmodes [9, 10, 11, 12, 13, 14, 15], either in their bulk or at their interface with air. This leads to exciting opportunities for nanoscience, with directional and highly confined optical beams supported by such hyperbolic materials. Their features can be accessed by exciting hyperbolic media with localized sources. Generally, localized excitations of a medium access all available momenta, leading to an emission pattern that is the spatial Fourier transform of the iso-frequency dispersion contour. For isotropic media, a localized emitter generates a field profile that follows a bullseye pattern, which is indeed the Fourier transform of a circular iso-frequency dispersion curve [16]. Similarly, if a material supports a hyperbolic iso-frequency contour at a given frequency, the emission from a localized emitter will feature a field distribution associated with the Fourier transform of a hyperbola. In order to image its pattern, we can rely on near-field scanning optical experiments, as done in many recent papers [17, 18, 19, 20], or rely on another platform to realize optically the Fourier transform of a spatial pattern, i.e., using an optical lens [21]. Figure 1a shows the experimental image plane distribution created by a lens, while imaging two slits with hyperbolic profile in the object plane. In the following, we analytically evaluate this field pattern, and apply this result to provide physical insights into wave propagation in hyperbolic media and other phenomena featuring the Fourier transform of a hyperbolic curve.
@@ III Graphical representation
We start with this simplest region, which is characterized by a pair of straight lines given by $y=\pm\sqrt{-\nicefrac{{\epsilon_{y}}}{{\epsilon_{x}}}}\,x$. The functional value of the transform for all points in this region is unbounded. This “blow-up” is acceptable since our hyperbola has an infinite support. Note that, as the two asymptotes of the given hyperbola in reciprocal space are described by $k_{y}=\pm\sqrt{-\nicefrac{{\epsilon_{x}}}{{\epsilon_{y}}}}\,k_{x}$ (see Fig. 2a which depicts the case of $\epsilon_{x}<0$ and $\epsilon_{y}>0$), each of the asymptotes is normal to one of the separatrix lines, as shown in Fig. 2b. This orthogonality relation can be understood in the limit of $\kappa\rightarrow 0$ where our hyperbola evolves into the degenerate case of a pair of straight lines. In this limiting case, the corresponding transform (see Eq. (2)) becomes zero in the major and minor regions, and blows up in the separatrix, giving us a pair of straight lines orthogonal to the original degenerate pair, which is expected based on our discussion of the Fourier transform of a straight line in Section II. Moreover, this $\kappa\rightarrow 0$ picture also explains why the description of the separatrix is unaltered if the signs of $\epsilon_{x}$ and $\epsilon_{y}$ are reversed. In fact, as the description does not depend on the value of $\kappa$ at all, and hence where the two foci of the hyperbola are located, one can conclude that the separatrix captures the Fourier transform information of the asymptotic behavior of a hyperbola.
@@
Combining the individual contributions of all three regions gives us the complete graphical representation of the Fourier transform of a hyperbola, akin to Fig. 1b, corresponding to the fields radiated by a point source in a hyperbolic medium. Note the coherence among the three regions with a blow-up in the separatrix and decay on either side. Interestingly, the decay in the major region is much slower than that in the minor region. This is due to the fact that, away from the origin, the transform value in the major region scales as $1/{\sqrt{\kappa\Lambda}}$, whereas in the minor region it scales as ${e^{-\kappa\Lambda}}/{\sqrt{\kappa\Lambda}}$ [23]. As the dispersion hyperbolas change in eccentricity, orientation or size, the corresponding transform patterns change accordingly, as shown in Fig. 5. For hyperbolas with smaller eccentricity, the major region expands (see Fig. 5a), whereas for larger eccentricity it contracts (see Fig. 5b). Interchanging the directions of major and minor axes likewise swaps the corresponding major and minor regions in the transform (see Fig. 5c). When the overall size of the hyperbola is scaled up, the fringes become more closely spaced (see Fig. 5d). Now, why is the Fourier transform of a hyperbola the way it is, why does it have fringes on one region while in the other region it is structureless, how do we interpret it physically – we answer these questions in the next section.
Figure 5: Variations in eccentricity, orientation and overall size of the dispersion hyperbola (left panels) modify the transform pattern (right panels). For hyperbolas described by (1), here we considered (a) $3\epsilon_{x}=\epsilon_{y}=9$, (b) $\tfrac{1}{3}\epsilon_{x}=\epsilon_{y}=9$, (c) $\epsilon_{x}=-\epsilon_{y}=9$, and (d) $\epsilon_{x}=-\epsilon_{y}=27$. In all cases, the exponential profile in the minor region is overshadowed by the much larger transform amplitudes in the major region.
@@ IV Physical Interpretation
This additional correspondence gives us the direction of group velocity of the wave for each sector of the field profile, and affirms that the energy indeed propagates away from the source in a radially outward manner – a fact not obvious for a field distribution that is not circularly symmetric as a bullseye pattern.
@@ V Huygens’ principle in hyperbolic media
Figure 11a also illustrates how Huygens’ principle applied to hyperbolic wavelets naturally leads to the concept of negative refraction [38, 39, 40, 41]. For its key importance in modern optics, we highlight this effect separately in Fig. 12. When light refracts into a hyperbolic medium from an isotropic dielectric through an interface along which the permittivity component of the hyperbolic material is positive, Huygens’ construction, as shown in Fig. 12a, allows light to refract with a negative angle, and hence, energy flows the “wrong” way along the tangential direction. This is to be contrasted with the conventional (positive) refraction picture in Fig. 12b involving two dielectrics with different refractive indices which was originally studied by Huygens [28] through his circular wavelets.
@@ VII Conclusion
The authors thank Viktoriia Rutckaia and J. Ryan Nolen for providing the images used in Figs. 1a and 14a, respectively, and acknowledge support for this work from the Simons Foundation.